OpenStudy (anonymous):

use the differential equation given by dx equals x times y divided by 3, y > 0. Complete the table of values x −1 −1 −1 0 0 0 1 1 1 y 1 2 3 1 2 3 1 2 3 dy, dx and also Find the particular solution y = f(x) to the given differential equation with the initial condition f(0) = 4. I don't need the solutions, I just need to know how to do them, any help is appreciated!

OpenStudy (amistre64):

well, your table should be dx = xy/3 but its hard to see how to place them is x a function of y or some other random variable?

OpenStudy (amistre64):

oh, it says y=f(x)

OpenStudy (amistre64):

ive got nothing without a clear picture of the problem ..

OpenStudy (anonymous):

So it's not just me, thats the exact question typed up, I feel like my teacher pulled this question from the internet and didn't really think about it, I don't know how to do this question. Here is a picture of the question with proper formatting.

OpenStudy (anonymous):

@amistre64 Any clue,I still don't understand how to do this question, It's the only one I have left.

OpenStudy (amistre64):

thats better ..... the pic that is dy/dx is defined by xy/3 they give you xy, multiply them together and divide by 3 to determine dy/dx

OpenStudy (amistre64):

dy/dx = xy/3 dy/dx = -1(1)/3 dy/dx = -1(2)/3 dy/dx = -1(3)/3 dy/dx = 0(1)/3 dy/dx = 0(2)/3 dy/dx = 0(3)/3 dy/dx = 1(1)/3 dy/dx = 1(2)/3 dy/dx = 1(3)/3

OpenStudy (amistre64):

y' = xy/3 is a sovlable diffy q on its own ... dunno what the table would be for

OpenStudy (amistre64):

dy/dx = xy/3, its seperable, so seperate it 3 dy/y = x dx integrate it 3 lny = 1/2 x^2 + K and solve for y y = Ce^(x^2 /6) now define C using the intial condition .... its a very basic process here

OpenStudy (anonymous):

Thank you, I was confused and you cleared it up